### v2

# FairSteer: Inference Time Debiasing for LLMs with Dynamic Activation Steering

Abstract
Large language models (LLMs) are prone to capturing biases from training corpus, leading to potential negative social impacts. Existing prompt-based debiasing methods exhibit instability due to their sensitivity to prompt changes, while fine-tuning-based techniques incur substantial computational overhead and catastrophic forgetting. In this paper, we propose FairSteer, a novel inference-time debiasing framework without requiring customized prompt design or model retraining. Motivated by the linear representation hypothesis, our preliminary investigation demonstrates that fairness-related features can be encoded into separable directions in the hidden activation space. FairSteer operates in three steps: biased activation detection, debiasing steering vector (DSV) computation, and dynamic activation steering. Specifically, it first trains a lightweight linear classifier to detect bias signatures in activations, and then computes DSVs as intervention directions derived from small contrastive prompt pairs. Subsequently, it performs debiasing by adjusting activations with DSVs in the inference stage. Comprehensive evaluation with six LLMs demonstrates the superiority of FairSteer across question-answering, counterfactual input evaluation and open-ended text generation tasks.
FairSteer: Inference Time Debiasing for LLMs with Dynamic
Activation Steering
Yichen Li1thanks: Equal contribution., Zhiting Fan [^1] footnotemark: 1, Ruizhe Chen1, Xiaotang Gai1, Luqi Gong3, Yan Zhang4, Zuozhu Liu1,2thanks: Correspondence author.
1Zhejiang University 2Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence 3Zhejiang Lab 4ByteDance {yichen.23, zuozhuliu}@intl.zju.edu.cn
1 Introduction
Figure 1: Overview of the FairSteer framework. FairSteer operates in three steps: (a) Biased Activation Detection: train linear classifiers to detect bias signature. (b) Debiasing Steering Vector Computation: compute the DSV by averaging the differences between activations corresponding to biased and unbiased prompts. (c) Dynamic Activation Steering: apply conditional interventions in the selected layer.
Large language models (LLMs) have demonstrated remarkable performance in various tasks (Radford et al., 2019; Mann et al., 2020; Kojima et al., 2022). However, they tend to inherit social biases from their training corpus, leading to behaviors that negatively impact underrepresented groups (Zhao et al., 2019; Hutchinson et al., 2020; Sheng et al., 2021; Navigli et al., 2023; Mei et al., 2023). This challenge thus requires effective debiasing techniques to better align LLMs with ethical AI principles.
Existing debiasing techniques focus primarily on two paradigms: (1) In-context prompting-based methods (Dong et al., 2023; Gallegos et al., 2024b; Oba et al., 2024; Sun et al., 2024) leverage carefully crafted instructions to guide LLMs toward unbiased outputs, yet their effectiveness relies heavily on the quality of the prompt and is not robust to variations in phrasing. (2) Fine-tuning-based methods (Zmigrod et al., 2019; Ravfogel et al., 2020; Zayed et al., 2024; Liu et al., 2024; He et al., 2022; Allam, 2024) typically retrain models using balanced datasets or apply methods such as projection-based techniques, component-specific debiasing, contrastive learning and reinforcement learning. They provide more direct control over biases, with the cost of high computational expense, risk of catastrophic forgetting, and dependence on large annotated datasets that are hard to collect.
Therefore, the question arises: can we mitigate bias in LLMs without retraining, instead deferring the debiasing process to the inference stage? This approach, which we categorize as inference-time debiasing, is simpler and more practical than in-training debiasing, as it avoids the need for complex training procedures or extensive computational resources. Recent research on inference-time debiasing focuses on modifying decoding strategies (Saunders et al., 2021; Sheng et al., 2020; Meade et al., 2023; Lu et al., 2020b; Yang et al., 2022; Schick et al., 2021) by suppressing biased tokens. However, such strategies usually reduce the diversity of outputs and degrade the performance of LLMs on non-debiasing tasks.
Our work is initially inspired by the linear representation hypothesis (Bolukbasi et al., 2016; Mikolov et al., 2013; Elhage et al., 2022; Park et al., 2023; Jiang et al., 2024): semantic features like truth (Li et al., 2024), sentiment (Tigges et al., 2023), humor (von Rütte et al., 2024), and refusal (Arditi et al., 2024) are encoded as linearly separable directions in LLM activation spaces. This raises a fundamental question: Can fairness-related concepts similarly be encoded as separable geometric structures in hidden states, enabling bias mitigation through activation steering? To investigate this, we perform a preliminary analysis across six LLMs, as shown in Figure 2. Our key observation reveals that bias signatures exhibit over 90% linear separability in intermediate layers, indicating the feasibility of geometric intervention.
Motivated by our preliminary findings, we propose FairSteer, a novel inference-time debiasing method that dynamically adjusts activation without retraining. Our approach operates in three synergistic stages, as shown in Figure 1. (1) Biased Activation Detection (BAD): We train linear classifiers (Li et al., 2024; Xu et al., 2024) on intermediate layer activations to detect bias signatures across LLM layers with a predefined bias classification dataset. (2) Debiasing Steering Vector (DSV) Computation: We compute geometrically interpretable intervention directions using contrastive prompt pairs (Panickssery et al., 2023; Zou et al., 2023), isolating bias-specific features by controlling contextual variables. DSVs require only one hundred annotated examples, significantly fewer than the datasets for fine-tuning-based debiasing methods. (3) Dynamic Activation Steering (DAS): during inference, DAS acts as a plug-in and is applied only when a bias was detected, thereby better preserving the model’s original capabilities.
We conduct comprehensive experiments over six LLMs (Llama, Vicuna, Mistral) and four popular datasets (BBQ (Parrish et al., 2021), UNQOVER (Li et al., 2020), CrowS-Pairs (Nangia et al., 2020) and CEB (Wang et al., 2024)), with tasks across question answering, counterfactual input evaluation, and open-ended text generation. The results demonstrate FairSteer’s superior debiasing performance across different evaluation metrics, while nearly preserving the original performance of LLMs on MMLU/ARC/OBQA knowledge tasks. Extensive analysis, ablation studies, and case studies further highlight the effectiveness of BAD and DAS in FairSteer, suggesting the great potential of inference-time debiasing. The code is available at https://github.com/LiYichen99/FairSteer.
2 Preliminary
2.1 Problem Formulation
Let a language model $\mathcal{M}$ with $L$ transformer layers process an input sequence $P$ with $n$ tokens, i.e., $P=(t_{1},\dots,t_{n})$. For the token $t_{i}$ at position $i$ in layer $l$, we denote its residual stream activation as $\mathbf{a}_{i}^{l}\in\mathbb{R}^{d}$, where $d$ is the hidden dimension. In this work, we focus on the last token’s activation (i.e., $\mathbf{a}_{n}^{l}$ corresponding to $t_{n}$ in layer $l$), and simplify its notation as $\mathbf{a}^{l}$ for layer $l$. Our goal is to mitigate biases in $\mathcal{M}$’s next-token prediction $t_{n+1}$ during inference by dynamically adjusting $\mathbf{a}^{l}$.
Figure 2: Accuracy on validation set across different layers of LLMs.
Figure 3: 2-D PCA visualization of biased and unbiased activations from the 14th layer of Vicuna-13B. The arrow represents the mean difference between biased and unbiased activations, corresponding to the DSV.
2.2 Hypothesis: Linear Separability of Bias
To validate the linear representation hypothesis for fairness-related features, we conduct two exploratory experiments.
Layer-wise Separability Analysis. We train the linear classifier $C^{l}$ to distinguish biased and unbiased activations $\mathbf{a}^{l}$ across all layers. As shown in Figure  2, the validation accuracy peaks at intermediate layers (e.g., exceeding 90% from layer 14 in Mistral-7B) and remains stable above 87% in the final layer for all tested LLMs, indicating that fairness concepts become linearly separable from intermediate layers.
Geometric Subspace Visualization. To further validate separability, we project the activations of layer-14 of Vicuna-13B into the 2D space by principal component analysis (PCA) in Figure 3. The projection reveals a clear separation between biased (red) and unbiased (green) clusters. The Debiasing Steering Vector (DSV), computed as the mean difference between clusters, effectively bridges these subspaces. This geometric property confirms that bias mitigation can be operated through vector space interventions.
3 Methodology
Based on our preliminary findings, we introduce FairSteer which operates in three stages: 1) Biased Activation Detection, 2) Debiasing Steering Vector Computation, and 3) Dynamic Activation Steering during inference, as shown in Figure 1.
3.1 Biased Activation Detection
To enable precise control over debiasing interventions while preserving model capabilities, we train lightweight linear classifiers to detect bias signatures in real time during inference, serving as triggers for conditional intervention. First, We construct the dataset $\mathcal{D}_{\text{BAD}}$, where we label the model’s response as biased ($y=0$) if it selects the stereotypical answer, and unbiased ($y=1$) if it provides a neutral answer. For each sample, we extract the last token’s activation $\mathbf{a}^{l}$ from each layer $l$, which encapsulates the model’s compositional reasoning state before generating. Next, for each layer $l$, we train a linear classifier $C^{l}$, where the predicted label $\hat{y}$ for a given activation $\mathbf{a}^{l}$ is computed as:
$\hat{y}=\sigma(\mathbf{w}^{T}\mathbf{a}^{l}_{i}+b)$ (1)
Here, $\sigma$ is the sigmoid activation function, $\mathbf{w}$ is the weight vector, and $b$ is the bias term. The classifier is trained using a cross-entropy loss function with regularization:
$\mathcal{L}$ $=-\frac{1}{|\mathcal{D}_{\text{BAD}}|}\sum_{(\mathbf{a}^{l},y)\in\mathcal{D}_{\text{BAD}}}\Big{[}y\log(\hat{y})$ (2)
$+(1-y)\log(1-\hat{y})\Big{]}+\lambda\|\mathbf{w}\|_{2}^{2}$
where $\lambda$ is the regularization parameter. Implementation details can be found at Appendix C.2.
3.2 Debiasing Steering Vector Computation
Figure 4: Contrast prompt pairs example.
Based on the validation in Section 2.2, we can mitigate bias through geometric intervention by computing a Debiasing Steering Vector (DSV), which captures the directional offset between biased and unbiased activation subspaces. To compute the DSV, we first construct a dataset $\mathcal{D}_{\text{DSV}}$ that contains $N$ contrast prompt pairs $(\mathcal{P^{+}},\mathcal{P^{-}})$: $\mathcal{P^{+}}$ demonstrates biased behavior, while $\mathcal{P^{-}}$ demonstrates unbiased behavior. As shown in Figure 4, $\mathcal{P^{+}}$ and $\mathcal{P^{-}}$ share identical contexts but differ in answer choices to elicit unbiased and biased responses, respectively. This design ensures that the DSV captures the desired fairness-related features while minimizing the influence of unrelated factors. Then, the DSV $\mathbf{v}^{l}$ for layer $l$ is computed by averaging the differences between activations corresponding to biased and unbiased prompts:
$\mathbf{v}^{l}=\frac{1}{|\mathcal{D}_{\text{DSV}}|}\sum\limits_{(\mathcal{P^{+}},\mathcal{P^{-}})\in\mathcal{D}_{\text{DSV}}}[\mathbf{a}^{l}(\mathcal{P^{+}})-\mathbf{a}^{l}(\mathcal{P^{-}})]$ (3)
Note that, the DSV encodes both directional and magnitude information: its direction represents the optimal debiasing trajectory from biased to unbiased subspaces, while its magnitude quantifies the average distance between these subspaces.
3.3 Dynamic Activation Steering
To balance debiasing efficacy with model capability preservation, FairSteer employs conditional interventions triggered only when biases are detected. This dynamic mechanism avoids distorting unbiased outputs while ensuring precise corrections for biased generations.
Given an input prompt $P$, we first extract the last token’s activation $\mathbf{a}^{l^{*}}(P)$ in the pre-selected layer $l^{*}$ (as detailed in Section 4.1). We then use the pre-trained classifier $C^{l^{*}}$ to compute the bias probability $\hat{y}=C^{l^{*}}(\mathbf{a}^{l^{*}}(P))$. The intervention is triggered if $\hat{y}<0.5$, indicating biased activation. Once triggered, we apply the DSV to adjust the activations as follows:
$\mathbf{a}^{l^{*}}_{\text{adj}}(P)=\mathbf{a}^{l^{*}}(P)+\mathbf{v}^{l^{*}}$ (4)
Here, the adjusted activation $\mathbf{a}^{l^{*}}_{\text{adj}}(P)$ propagates through subsequent layers, steering the generation toward unbiased outputs.
4 Experiments
Figure 5: Layer selection based on accuracy across six models.
Question-Answering Counterfact Text Generation
BBQ UNQOVER CrowS CEB
ZS FS ZS FS
[…]
Llama2-7B 32.21 2.44 2.41 42.67 6.43 5.36 7.03 20.60 67.75 0.68 0.0119 0.54 21.24
CAL 41.40 0.87 1.62 38.75 6.23 6.99 62.52 12.30 - - - - -
Ours 46.28 1.04 2.11 53.34 2.63 4.55 19.67 28.52 66.51 0.69 0.0144 0.62 20.42
Llama2-13B 48.60 5.86 2.91 47.94 16.31 5.55 33.96 19.17 70.93 0.69 0.0105 0.51 23.44
CAL 51.29 1.41 2.46 53.27 9.82 5.47 60.32 32.27 - - - - -
Ours 74.02 -0.82 0.84 80.26 1.58 3.68 53.00 49.23 69.46 0.75 0.0081 0.72 19.81
Llama3-8B 71.00 13.62 2.51 84.74 13.53 2.42 20.84 76.22 67.83 0.76 0.0078 0.61 19.38
CAL 55.51 0.08 5.64 82.65 2.61 2.69 99.75 95.67 - - - - -
Ours 90.22 1.46 2.17 92.12 4.39 2.32 58.01 91.94 66.82 0.70 0.0071 0.65 19.22
Vicuna-7B 41.33 6.78 5.94 43.89 14.28 9.07 16.19 18.34 69.53 0.66 0.0178 0.69 17.74
CAL 33.45 -0.01 -0.02 40.34 15.33 9.73 33.66 10.88 - - - - -
Ours 65.38 1.47 5.16 71.28 2.80 7.80 43.11 57.21 68.06 0.67 0.0123 0.84 15.51
Vicuna-13B 63.71 4.97 3.56 64.74 15.72 5.49 41.44 52.90 69.92 0.72 0.0131 0.69 20.31
CAL 47.99 0.72 1.23 63.72 12.11 5.93 35.34 58.76 - - - - -
Ours 77.74 0.10 2.50 86.56 1.28 4.33 49.06 73.19 69.30 0.80 0.0055 0.82 14.19
Mistral-7B 62.22 10.30 5.24 70.53 16.68 5.10 33.06 57.83 71.47 0.66 0.0148 0.53 22.84
CAL 68.32 2.28 4.00 81.89 8.43 4.68 67.54 87.02 - - - - -
Ours 73.43 3.75 5.02 79.93 8.04 4.87 54.10 69.51 71.09 0.67 0.0184 0.54 21.87
Table 1: Comparison of debiasing performance between our method and baselines on three tasks. ZS and FS refer to zero-shot and few-shot settings. We use accuracy (Acc), bias score on ambiguous (BS(a)) and disambiguated (BS(d)) contexts, stereotype score (SS), sentiment (Senti), toxicity (Toxic), regard (Regard), and bias score (BS) as our metrics for evaluation. The best result is indicated in bold.
Figure 6: Accuracy across different categories of BBQ for Llama2-13B, comparing the Base Model, CAL, and Ours. Full results are provided in Appendix F.
4.1 Settings and Details
We conduct our experiments on multiple popular contemporary LLMs: Llama-2-chat (Touvron et al., 2023b) 7B and 13B (parameters), Llama-3-instruct (AI@Meta, 2024) 8B, Vicuna-v1.5 (Chiang et al., 2023) 7B and 13B, Mistral-v0.3-instruct (Jiang et al., 2023) 7B. For decoding, we employ greedy decoding in all experiments to ensure reproducibility. In the open-ended text generation task, we set the maximum length for new tokens to 512.
Layer Selection. To identify the optimal layer, we generate a dataset containing 2200 examples (200 per category) from BBQ. We evaluate the accuracy of each layer on this dataset and select the one with the highest accuracy as the optimal layer. As shown in Figure 5, we observe that the optimal layer for all LLMs tested lies between layer 13 and layer 15, within the intermediate layers. This aligns with the results in Figure 2, where the classifier also achieves peak accuracy in the intermediate layers. Additionally, this finding supports conclusions from (Skean et al., 2024), which suggest that intermediate layers often yield more informative and generalizable features, as they strike a balance between low-level token representations and high-level semantic features. These layers efficiently compress and refine information, isolating the most relevant features for tasks such as bias detection, while preserving a robust representation of both fine-grained and abstract knowledge. Therefore, intermediate layers are the optimal layer for our method, supporting stable and effective bias mitigation without compromising performance.
4.2 Datasets
Datasets for Biased Activation Detection. We construct the dataset by combining 58,492 examples from BBQ and 10,266 examples from MMLU (Hendrycks et al., 2020). This mixture prevents classifier overfitting to domain-specific artifacts while maintaining discrimination capability across different bias categories. The dataset is split into training and validation sets in a 4:1 ratio, and the classifier is fitted on the training set. Further details are provided in Appendix C.1.
Datasets for DSV Computation. We construct contrast prompt pairs from BBQ. We select BBQ due to its broad coverage of social biases, including nine categories and two intersectional biases. We sample 10 examples from each category, collecting a total of 110 examples to serve as the dataset for DSV computation. Further details and the impact of dataset size are discussed in Appendix D.
4.3 Evaluation Tasks
We evaluate debiasing performance on three tasks:
Question-Answering. We conduct our experiments on BBQ and UNQOVER. BBQ contains 58,492 questions across nine categories, while UNQOVER includes 40,000 questions across four categories. In both datasets, we use accuracy as the evaluation metric. Additionally, for BBQ, to quantify the extent to which a model systematically provides biased responses, we calculate bias scores separately for ambiguous and disambiguated contexts as defined by Parrish et al. (2021). A detailed description of the bias score calculation is provided in Appendix E.1.
Counterfactual input evaluation. We use CrowS-Pairs containing 1,508 examples across nine types of bias. We measure the model’s preference for the stereotypical sentence by calculating the perplexity of both sentences in a zero-shot setting, with higher scores indicating a stronger bias.
Open-ended text generation. We use CEB and focus on the continuation task related to stereotyping. This dataset is derived from the HolisticBias (Smith et al., 2022) and includes 400 examples across four social groups. We use different bias metrics provided by Dhamala et al. (2021) and Wang et al. (2024) to capture biases from multiple perspectives: (1) Sentiment. We use the VADER (Hutto and Gilbert, 2014) sentiment analyzer to evaluate the sentiments conveyed in the texts generated by LLMs. (2) Toxicity. We use unitaryai/detoxify [^1] https://github.com/unitaryai/detoxify library to measure the toxicity. (3) Regard. Sheng et al. (2019) introduced regard as an alternative to sentiment, measuring the social perception of a demographic in text. Contrary to toxicity, higher scores for regard and sentiment indicate positive connotation. (4) Bias score. we use GPT-4 (Achiam et al., 2023) to assess the degree of stereotyping in the generated outputs, with scores ranging from 0 (nonbiased) to 100 (most biased). Specific prompts from Wang et al. (2024) are provided to GPT-4 to guide the scoring process, as detailed in Appendix E.2.
4.4 Baseline Methods
We compare our methods with several baseline approaches: Zero-shot and few-shot baselines are used for evaluation, with few-shot prompts from Si et al. (2022). Causal-Guided Active Learning (CAL) (Sun et al., 2024) leverages the model’s capabilities to identify biased samples and patterns, then applies in-context learning method to prevent bias during generation. In the zero-shot setting, CAL induces bias patterns from the data and appends a debiasing prompt to the original prompt. In the few-shot setting, CAL creates counterfactual examples to guide the model away from biased patterns. Note that, since Crows-Pairs and CEB-continuation datasets are not suitable for few-shot scenarios and lack a clear set of possible answers, we are unable to provide counterexample pairs for CAL to induce bias pattern. Therefore, we only test the baselines on question-answering tasks.
4.5 Main Results
The main results are presented in Table 1, where we compare our method with baselines across six models and three tasks:
Question-Answering on BBQ and UNQOVER. First, compared to the base model, our method effectively improves accuracy and reduces bias scores in both ambiguous and disambiguated contexts across all tested models, in both zero-shot and few-shot settings. This demonstrates the effectiveness and robustness of our approach in reducing bias in question-answering tasks. Second, in the BBQ results, FairSteer consistently outperforms CAL in terms of accuracy. Although CAL reduces bias scores for some models, it does so at the expense of accuracy. For instance, in the Vicuna-7B zero-shot setting, CAL achieves a significantly low bias score but results in lower accuracy compared to the base model. Third, in the UNQOVER results, while some models in the zero-shot setting perform worse than CAL, FairSteer outperforms CAL in most few-shot settings. Moreover, we observe that few-shot FairSteer consistently outperforms zero-shot FairSteer in accuracy, as the inclusion of additional correct instances in few-shot settings assists in guiding the model toward more accurate answers. In contrast, CAL exhibits inconsistent performance across models. In the case of Llama2-7B, Vicuna-7B and Vicuna-13B, adding CAL even results in a performance decline compared to the base model, indicating that CAL does not consistently improve bias mitigation and may even negatively affect model accuracy. Overall, FairSteer demonstrates more stable debiasing performance across different models and tasks, proving to be more reliable and effective than CAL in both zero-shot and few-shot settings.
Counterfactual input evaluation on CrowS-Pairs. Our method reduces the stereotype score across all tested models, demonstrating its effectiveness in decreasing the model’s preference for stereotypical sentences and mitigating social bias.
Open-ended text generation on CEB. Our method shows promising results, with consistent reductions in bias score across all tested models. Additionally, improvements in sentiment, toxicity, and regard metrics further validate its effectiveness in reducing bias while maintaining the quality of the generated content.
Figure 7: Case study examples. Further examples are provided in Appendix G.
4.6 Analysis and Discussion
Results across Different Categories. BBQ consists of nine categories and two intersectional biases. In Figure 6, we compare the accuracy and bias scores of FairSteer and baselines across all categories in both zero-shot and few-shot settings. Our results show that FairSteer improves accuracy across all categories and reduces bias scores in nearly all of them. Compared to CAL, FairSteer outperforms it in terms of accuracy in all categories and achieves lower bias scores in most. No single category appears to be solely responsible for the overall performance, and no clear pattern emerges as to which categories exhibit the most significant effects. Additionally, we analyze the impact of different categories on BAD and DSV in Appendix A.
Impact on General Model Performance. We conduct experiments on two types of tasks: language modeling and general downstream tasks. For language modeling, we measure the perplexity on the WikiText 103 corpus (Merity et al., 2016), using HuggingFace’s evaluate library. [^2] https://github.com/huggingface/evaluate For general downstream tasks, we evaluate on three datasets: (1) MMLU, contains 14,042 questions across 57 tasks, offering broad subject coverage and varying levels of difficulty. (2) ARC (Clark et al., 2018), consists of 7,787 grade-school science questions, divided into Challenge and Easy Sets. (3) OpenBookQA (OBQA) (Mihaylov et al., 2018), consists of 500 questions that evaluate the model’s ability to use provided scientific facts to answer related questions. We use a 5-shot setting for MMLU and zero-shot for ARC and OBQA, following the methodology in Touvron et al. (2023a).
In Table 2, the results demonstrate that our method has minimal impact on general task performance, with only a slight reduction in language modeling. Accuracy on MMLU, ARC, and OBQA remains mostly unchanged, suggesting that our method effectively reduces bias without compromising overall performance.
LM General Downstream
[…]
Llama2-7B 31.75 47.13 74.45 56.40 58.80
Ours 31.52 46.73 74.45 56.31 58.80
Llama2-13B 31.29 53.53 81.36 66.72 64.00
Ours 31.61 53.35 81.27 66.64 63.80
Llama3-8B 188.21 68.37 93.56 83.53 81.60
Ours 188.69 68.34 93.56 83.53 81.60
Vicuna-7B 34.73 49.90 75.59 57.08 60.60
Ours 36.63 49.59 75.59 57.08 60.60
Vicuna-13B 15.37 55.88 83.25 68.26 64.40
Ours 15.53 55.76 83.25 68.26 64.40
Mistral-7B 36.84 61.94 88.05 77.05 75.60
Ours 36.57 61.95 88.05 77.05 75.40
Table 2: Performance comparison of capability on language modeling and general downstream tasks.
Zero-Shot Few-Shot
[…]
Llama2-7B 32.21 2.44 2.41 42.67 6.43 5.36
DSV 39.30 0.93 1.07 45.25 2.16 4.13
Ours 46.28 1.04 2.11 53.34 2.63 4.55
Llama2-13B 48.60 5.86 2.91 47.94 16.31 5.55
DSV 52.84 0.05 -1.15 55.46 1.96 1.77
Ours 74.02 -0.82 0.84 80.26 1.58 3.68
Llama3-8B 71.00 13.62 2.51 84.74 13.53 2.42
DSV 62.21 0.71 3.09 74.11 3.42 3.83
Ours 90.22 1.46 2.17 92.12 4.39 2.32
Vicuna-7B 41.33 6.78 5.94 43.89 14.28 9.07
DSV 55.48 0.29 1.86 55.66 2.49 9.03
Ours 65.38 1.47 5.16 71.28 2.80 7.80
Vicuna-13B 63.71 4.97 3.56 64.74 15.72 5.49
DSV 55.48 0.29 1.86 61.43 1.03 5.89
Ours 77.74 0.10 2.50 86.56 1.28 4.33
Mistral-7B 62.22 10.30 5.24 70.53 16.68 5.10
DSV 59.76 2.29 4.91 74.23 6.42 5.45
Ours 73.43 3.75 5.02 79.93 8.04 4.87
Table 3: Ablation study results on BBQ. DSV refers to FairSteer with the BAD removed.
Ablation Study. To investigate the role of BAD, we remove it and use only DSV for inference guidance, testing the results on BBQ, as shown in Table 3. In both zero-shot and few-shot settings, we find that removing BAD and using only DSV still improves accuracy and reduces bias scores in most models, indicating that DSV alone can achieve some bias mitigation. However, using only DSV results in a significant drop in accuracy for most models compared to our full method. While DSV does reduce bias scores, the accuracy loss highlights the importance of BAD in guiding the model toward more accurate responses while maintaining bias mitigation. This indicates that BAD helps strike a better balance between accuracy and bias reduction, leading to more stable and reliable results. Thus, while DSV alone has some bias mitigation capability, the full method with BAD outperforms DSV in both accuracy and bias mitigation across all tested models.
Case Study. To clearly demonstrate the effectiveness of our method, we present two case studies. Figure 7(a) shows an example from question-answering, where Llama2-13B generates the biased answer “B. The grandson”. Our method adjusts the output to “C. Can’t be determined”, highlighting the ambiguity of the passage. Moreover, our method provides a logically coherent explanation, further supporting the conclusion that the correct answer cannot be definitively determined from the given information. This example shows how our method mitigates bias by correcting responses that might otherwise reflect stereotypical assumptions.
Figure 7(b) illustrates an example from open-ended text generation, where Llama2-13B generates a harmful statement about trans masc individuals. Our method generates a more neutral and inclusive response, reframing the language to promote respect and dignity for all individuals, regardless of gender identity. This example highlights how our method reduces toxicity while maintaining overall coherence and relevance.
5 Related Work
Promoting fairness in LLMs is a critical component of alignment, a broader objective that seeks to ensure AI systems adhere to human values Stiennon et al. (2020); Ouyang et al. (2022); Fan et al. (2024a); Chen et al. (2024d, a, 2025). Researchers have proposed various debiasing methods, which can be group into two main paradigms: (1) In-context prompt-based methods (Dong et al., 2023; Gallegos et al., 2024b; Oba et al., 2024; Sun et al., 2024) leverage carefully crafted instructions or contextual prompts to guide LLMs toward unbiased outputs during generation. These approaches avoid modifying model parameters and are easily applicable to black-box models. However, they are sensitive to prompt formulation, struggle with implicit biases, and depends heavily on prompt quality and specific use cases. (2) Fine-tuning-based methods involve constructing rebalanced datasets and retraining models (Lu et al., 2020a; Webster et al., 2020; Zmigrod et al., 2019; Maudslay et al., 2019; Zayed et al., 2023; Chen et al., 2024c), or using techniques such as projection-based methods (Bolukbasi et al., 2016; Ravfogel et al., 2020; Liang et al., 2020), debiasing specific components (Gaci et al., 2022; Zayed et al., 2024; Liu et al., 2024; Limisiewicz et al., 2023; Chen et al., 2024b), contrastive learning (Cheng et al., 2021; He et al., 2022; Oh et al., 2022), adversarial learning (Han et al., 2021; Jin et al., 2020), and reinforcement learning (Liu et al., 2021b; Allam, 2024) approaches. Although effective, these methods involve complex and resource-intensive training, requiring diverse debiasing datasets.
In contrast, inference-time debiasing methods are simpler and more practical, as they do not require retraining or altering the model’s architecture. Existing inference-time techniques primarily focus on decoding strategy modification, such as constrained next-token search (Saunders et al., 2021; Sheng et al., 2020; Meade et al., 2023; Lu et al., 2020b) or generating and reranking alternative outputs (Chung et al., 2023; Kim et al., 2022; Liu et al., 2023, 2021a; Hallinan et al., 2022; Fan et al., 2024b). However, the key challenge with these methods is balancing effective bias mitigation while preserving output diversity (Gallegos et al., 2024a). Furthermore, many of these techniques (Yang et al., 2022; Schick et al., 2021) are tailored to older models such as BERT (Devlin et al., 2019) and GPT-2 (Radford et al., 2019), limiting their applicability to more recent architectures. Our approach advances inference-time debiasing by demonstrating that fairness-related features are linearly separable in activation space, enabling targeted interventions via geometrically interpretable steering vectors.
6 Conclusion
In this paper, we propose FairSteer, an inference-time debiasing framework for LLMs without requiring retraining. Inspired by the linear representation hypothesis, our preliminary analysis demonstrates that fairness-related features can be encoded into separable directions in the hidden activation space. FairSteer detects bias signatures in these activations using a lightweight linear classifier and applies debiasing steering vectors, which are computed by a small set of contrast prompt pairs, to dynamically adjust these activations during inference. Experimental results show that FairSteer effectively mitigates bias in tasks including question-answering, counterfactual input evaluation, and open-ended text generation, showcasing its broad applicability while preserving language modeling capabilities.
Limitations
Although our study provides valuable contributions, it has several limitations. First, the reliance on a linear classifier may limit its ability to detect more complex, non-linear forms of bias embedded in model activations. Second, the effectiveness of the debiasing steering vector depends on the quality and representativeness of the biased-unbiased prompt pairs used during its construction, which may not capture all types of bias present in real-world scenarios. Third, the approach we employ to derive the debiasing steering vector may not be optimal. This work serves primarily as a proof of fairness concept, demonstrating the existence of such a debiasing direction, rather than a thorough exploration of the most effective extraction techniques. Future research is needed to refine and enhance this methodology. Finally, while we assess performance across six open-source models, the generalizability of our findings to other models, especially large-scale, state-of-the-art proprietary systems and future architectures remains uncertain.
Acknowledgement
This work is supported by the National Natural Science Foundation of China (Grant No. 62476241), the Natural Science Foundation of Zhejiang Province, China (Grant No. LZ23F020008), Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence, and the Zhejiang University-Angelalign Inc. R&D Center for Intelligent Healthcare.
Appendix A Social Biases across Different Social Groups
Social biases can be categorized by social group, and different social groups may reflect different biases and fairness-related features. Therefore, we explore the impact of different social groups on the Biased Activation Detection and DSV computation. In this section, we focus on the categories defined by the BBQ dataset, which includes nine categories and two intersectional biases. Additionally, we use Llama2-13B to conduct these exploratory experiments.
A.1 Impact of Social Groups on Biased Activation Detection
For each category’s BAD dataset, we split it into training and validation sets with a 4:1 ratio, following the same procedure used for training classifiers across all categories in the main text. Table 4 shows the number of samples in the training and validation sets for each category. Each category’s classifier is trained on its respective training set.
Category Train Val
Age 4580 1145
Disability_status 1833 459
Gender_identity 7168 1792
Nationality 3841 961
Physical_appearance 2006 502
Race_ethnicity 8279 2070
Race_x_SES 13721 3431
Race_x_gender 18707 4677
Religion 1515 379
SES 8385 2097
Sexual_orientation 1012 253
All 79263 19816
Table 4: Training and validation set sizes for each category in the BBQ dataset. The last row “All” refers to training classifiers across all categories in the main text, and the dataset for training this classifier also includes MMLU.
Figure 8 shows the accuracy variation across layers for each category, where we observe that all classifiers, trained on different categories, follow a similar upward trend as the layers increase, reaching higher accuracy with deeper layers and peaking at intermediate layers. This trend validates the linear representation hypothesis, indicating that the activations become increasingly linearly separable with deeper layers, particularly at the intermediate layers. We also observe that classifiers trained on different categories reach different peak accuracies. For instance, categories like Age and Sexual orientation exhibit higher accuracy, while others, including Disability status, Physical appearance, Religion, and Sexual orientation, show a more gradual improvement. Furthermore, we find that the categories with lower classifier accuracy correspond to those with the smallest training set sizes. Based on this observation, we infer that the size of the training set may be an important factor influencing BAD performance.
Figure 8: Accuracy variation across layers for classifiers trained on different categories of BBQ.
As shown in Figure 10, We evaluate each category’s classifier on the validation sets of all categories. The heatmap reveals how well each classifier performs on the different categories, with higher accuracy indicated by deeper red shades. We observe that each classifier achieves high accuracy (greater than 90) on its own category’s validation set but performs with lower accuracy (below 80) on validation sets from other categories. This variation underscores the importance of category-specific training data in improving performance across different social group categories. Moreover, the “All” row at the bottom shows consistently high accuracy (greater than 90) across all validation sets, indicating that training on all categories leads to robust generalization.
A.2 Impact of Social Groups on DSV Computation
For each category, we sample 100 examples from BBQ and compute the corresponding DSV. In Figure 11, we compare the similarity of DSVs across different categories. The heatmap visualizes the pairwise cosine similarity between the DSVs of each category, with values closer to 1 indicating higher similarity. We find that the similarity between DSVs of different categories exceeds 0.90, suggesting that fairness-related features across these categories are similarly encoded in the model’s activation space. Additionally, the DSV computed from all categories containing 110 examples exhibits similarity values greater than 0.95 when compared to the DSV of any other category. This demonstrates that the DSVs capture a common underlying structure in the model’s activation space, reinforcing the hypothesis that fairness-related features are uniformly represented across different social categories.
$\alpha$ Acc BS(a) BS(d)
0 48.60 5.86 2.91
0.1 50.65 4.69 2.72
0.5 64.98 0.96 1.70
1 74.02 -0.82 0.84
1.5 73.99 -0.85 0.82
2 70.87 -0.87 0.81
Table 5: Effect of varying intervention strength $\alpha$ on Llama2-13B on BBQ.
Appendix B Effect of Varying Intervention Strength
In our method, we apply a fixed intervention strength $\alpha$ of DSV, when the bias is detected ($y<0.5$). While simple, this heuristic does not account for the magnitude of bias in activations, which may lead to mildly biased activations and over-steering risk.
Therefore, we conduct an experiment to evaluate the effect of varying (intervention strength) on Llama2-13B’s performance for the BBQ task. As shown in Table 5, increasing marginally reduces BS(a) further but slightly degrades accuracy, suggesting diminishing returns and minor over-steering. While $\alpha=1$ works well globally, we find that local over-steering might occur for mildly biased activations. To mitigate this, we propose an adaptive scaling of the DSV based on the severity of the bias in the future work. Specifically, we can use the classifier’s confidence score $p(y=0)$ to modulate $\alpha$: reduce the intervention strength for borderline cases ($p(y=0)\approx 0.5$), and apply stronger corrections for instances of clear bias ($p(y=0)\gg 0.5$).
Appendix C Details of Biased Activation Detection
C.1 Datasets
We construst the dataset from BBQ and MMLU. We note that BBQ has only three options, whereas MMLU has four. Since the linear classifier is sensitive to simple selection features, to prevent the inclusion of option D from interfering with classifier training, we filter out MMLU examples where the correct answer is not D. Additionally, we remove option D when posing prompts. Figure 9 provides examples used for training the linear classifier. For each question in BBQ, we label model responses as biased if the stereotypical answer is selected, and unbiased if the correct answer is chosen. In contrast, for each question in MMLU, we label all model responses as unbiased to avoid overfitting to bias-only patterns, while preserving the model’s general knowledge representations during the debiasing process. Furthermore, to ensure that BAD is applicable to both zero-shot and few-shot settings, we augment the BBQ questions by incorporating a few-shot scenario for data collection.
Figure 9: Examples for Biased Activation Detection on BBQ and MMLU.
C.2 Implemental Details
When training linear classifiers, we use the default settings provided by the cuML [^3] https://github.com/rapidsai/cuml library. Specifically, we employ the cuml.linear _model.LogisticRegression implementation, which utilizes cross-entropy loss and L2 penalty regularization with a regularization strength of $\lambda=1$.
Appendix D Details of DSV Computation
D.1 Datasets
We construct prompt pairs from BBQ. Each question in BBQ provides a context, two social groups (with one of the groups being assigned a negative stereotype in that context), and three answer options: a target answer (the group that reflects the stereotype), an unknown answer (e.g. cannot be determined) and a non-target answer (the remaining group). The questions are divided into two types: ambiguous contexts that missing information necessary to answer the questions, and disambiguated contexts that provides the necessary information. Given that models tend to strongly rely on social biases when the context is ambiguous (Gallegos et al., 2024b), we use this subset to generate our steering vectors. We sample 10 examples from each category, collecting a total of 110 examples to serve as the dataset for DSV computation.
D.2 Impact of Dataset Size and Random Seed on DSV
$\boldsymbol{n}$ Accuracy
- 44.86
10 65.36
20 65.27
30 65.50
50 66.00
100 65.40
Seed Accuracy
- 44.86
0 65.41
42 65.36
123 66.54
999 65.23
1234 65.63
Table 6: Accuracy variations of Llama2-13B when extracting the DSV under different dataset sizes (left) and random seeds (right). The first row with ’-’ refers to the original model without applying our method.
To investigate the impact of dataset size and random seed on DSV extraction from BBQ, we test on the dataset designed for layer selection. Since BBQ consists of nine categories and two intersectional biases, let $n$ denote the size per category. In the dataset size experiment, we fix the random seed at 42 and incrementally increase $n$ to compute the DSV. For the random seed experiment, we fix $n$ at 10 and select five commonly used seeds. Table 6 presents the accuracy variations observed in these experiments. The results show that our method significantly improves the original accuracy of Llama2-13B and maintains stable accuracy across different dataset sizes and random seeds, demonstrating the robustness of our approach.
Appendix E Evaluation Details
E.1 Bias Score Metrics in BBQ
To quantify the extent to which a model systematically provides biased responses, we calculate bias scores separately for ambiguous and disambiguated contexts as defined by Parrish et al. (2021). These scores measure the frequency with which the model generates the biased target answer. A bias score of 0% indicates that no bias is detected, while a score of 100% signifies that all responses align with the targeted social bias, and -100% indicates that all responses oppose the bias.
The bias score in disambiguated contexts ($s_{\text{DIS}}$) is calculated as follows:
$s_{\text{DIS}}=2\big{(}\frac{n_{\text{biased\_ans}}}{n_{\text{non-UNKNOWN\_outputs}}}\big{)}-1$ (5)
Here, $n$ represents the number of examples in each response group, with $n_{\text{biased\_ans}}$ being the number of outputs reflecting the targeted social bias, and $n_{\text{non-UNKNOWN\_outputs}}$ being the total number of outputs that are not marked as “UNKNOWN” (i.e., including both target and non-target answers).
The bias score in ambiguous contexts ($s_{\text{AMB}}$) is calculated as follows:
$s_{\text{AMB}}=(1-\text{accuracy})s_{\text{DIS}}$ (6)
We scale the bias scores in ambiguous contexts by accuracy to account for the fact that a biased answer becomes more harmful when it occurs more frequently. This scaling is not necessary in disambiguated contexts, as the bias score is not solely determined by incorrect answers.
E.2 Prompt for Scoring on CEB
We evaluate model bias in text generation using the CEB (Wang et al., 2024) dataset. Following the evaluation metrics outlined by Wang et al. (2024), given an LLM-generated output, we use GPT-4 (Achiam et al., 2023) to obtain a bias score. Specifically, the score is set between 0 and 99, with detailed descriptions of the bias degrees corresponding to different intervals. The prompt from Wang et al. (2024) is shown in Figure 12.
Figure 10: Heatmap of classifier accuracy across all categories. Each cell represents the accuracy of a classifier trained on one category and evaluated on the validation set of another category.
Figure 11: Pairwise cosine similarity between the Debiasing Steering Vectors (DSVs) of different categories
Figure 12: Prompt used for scoring on CEB.
Appendix F Full Results for Llama2-13B on BBQ
In Figure 13 and Figure 14, we present the bias score results seperately on ambiguous and disambiguated contexts for Llama2-13B on BBQ.
Figure 13: Bias score on ambiguous contexts across different categories of BBQ for Llama2-13B, comparing the Base Model, CAL, and Ours.
Figure 14: Bias score on disambiguated contexts across different categories of BBQ for Llama2-13B, comparing the Base Model, CAL, and Ours.
Appendix G Additional Case Studies
See Figure 15.
Figure 15: Additional case studies on BBQ and CEB.